# Flexible Process Variant Binding in Information Systems with Software Product Line Engineering

### v1

@@ Abstract
Different organisations often run similar digitised business processes to achieve their business goals. However, organisations often need to slightly adapt the business processes implemented in an information system in order to adopt them. Various approaches have been proposed to manage variants in process models. While these approaches mainly deal with control flow variability, in previous work we introduced an approach to manage implementation variants of digitised business processes. In this context Software Product Line (SPL) Engineering was applied to manage a set of common core artefacts including a process model from which Process-Aware Information Systems (PAIS) can be derived, which differ in the implementation of their process activities. When deriving a PAIS, implementations are selected for each process activity and then included in the PAIS at compilation time. One challenge that has not yet been solved is giving users of digitised business processes the option of selecting several features at runtime, i.e. selecting multiple activity implementations at runtime. This paper extends our previous work by not only allowing for the selection of activity implementations at compile time, but also at start time and runtime. Consequently, it becomes possible to defer the decision as to which features should be selected to start time and runtime. Furthermore, multiple implementations of a particular activity may be selected and executed concurrently. As another challenge different organisation may want to collect and base their decision on different information in a digitised business process. Consequently, the presented approach also allows customising the input and output data of activities when deriving a PAIS for a specific organisation.
@@
Business Process Management Software Product Line Engineering Process Configuration Process Variability Software Reuse
@@ 1 Introduction
The right to self-administration of German municipalities leads to various variants of the same business process. For example, craftspersons may apply for a special parking permit, which allows them to park in areas in which regular citizen have to pay or in which generally parking is not permitted. Special parking permits for craftspersons are available, inter alia, in the German cities of Munich [^1] http://www.muenchen.de/dienstleistungsfinder/muenchen/1072021/, Constance [^2] https://service.konstanz.de/Leistung-Detailansicht/?id=6000803, and Stuttgart [^3] https://www.stuttgart.de/vv/leistungen/sonderparkausweise-fuer-gewerbetreibende-und-soziale-dienste.php. The business process for checking an application for the special parking permit is similar between the municipalities. However, there are slight differences among the municipalities. In some municipalities, the special parking permit is issued automatically, whereas in others a municipal employee needs to issue the parking permit manually. Other researchers have observed and investigated business process variability in municipalities as well [33], focusing on control flow variability.
Approaches have been proposed to deal with business process variability. Reference processes [52] and base processes [21], respectively, were introduced that may be configured to meet the needs of an individual organisation. However, these approaches focus on variations in the control flow of a business process rather than on the implementation level. Software Product Line (SPL) Engineering [6] addresses the challenge of developing and maintaining a set of similar software products. A Software Product Line comprises a common set of core artefacts from which features may be selected in order to build a specific software product. In literature, the process of selecting features from an SPL and building a software product is referred to as product derivation [19][28]. A Process-Aware Information System (PAIS) corresponds to a software product that executes a business process involving human actors, applications, and information sources [20]. In previous work [24], we applied and simplified the concepts of SPL Engineering to PAIS (PAIS Product Line). At build time, any implementation can be selected for an activity. However, this static approach lacks flexibility, which is of utmost importance in practice.
@@ 1.1 Problem Statement
Furthermore, the implementations for the activities need to be selected at build time. Selection at start time or runtime is not possible. In order to give the process engineer of a business process the option of selecting features, the decision as to which activity implementations are selected must be deferred to startup time or runtime. In previous work, we used the build tool Apache Maven to conditionally package the selected activity implementations. When activity implementations are to be selected at startup time or runtime a build tool cannot be used anymore. The challenge is to find and adopt a mechanism from SPL Engineering allowing for the selection of activity implementations at startup time and runtime. However, note that when it becomes possible to select and, consequently, deselect activity implementations at runtime, it needs to be specified what shall happen with the data of a running implementation that gets deselected.
@@ 1.2 Contribution
This work addresses the aforementioned limitations and problems of PAIS Product Lines. The main contribution of this work is threefold:
@@
Moreover, it is assured that the process of designing, developing, selecting activity implementations, and deriving products (known as the phases of SPL Engineering) is tool supported.
@@ 1.3 Outline
The remainder of this paper is structured as follows: Section 2 presents basic concepts of SPL Engineering, which lay the foundation of the approach and which are necessary for understanding it. The approach of PAIS Product Lines we introduced in [24] is described in Section 3. In Section 4, this approach is extended to enable more flexibility in that the data structure may be customised during PAIS Product derivation, multiple implementation for one activity may be selected, and the implementation may be bound at different times. The flexibility of the presented approach is discussed in Section 6. Section 7 discusses related work. Section 8 concludes the paper and provides and outlook on open challenges and further research topics.
@@ 2 Fundamentals
In this section, basic concepts of SPL Engineering are presented, which constitute the foundation of our approach. The concepts deal with variability among similar software products that may be derived from a set of common core artefacts.
2.1 Variability Modelling
Kang et al. [27] introduced a notation to capture the functionality of a software product from the perspective of a user. By applying this notation, it becomes possible to outline the features of a software product as well as their relationship in a tree-structured feature model. The root node represents the software product. Features of the software product are connected by edges to the root node. Features themselves may consists of other features and are inter-connected with them via edges as well. Little circles at the end of an edge refer to optional features, whereas alternative features are identified by arcs between their corresponding edges. Feature diagrams help developers to identify what has to be configured in a software product.
An example of a feature model is displayed in Figure 1a. Software Product 1 consists of three feature, i.e. Features f1, f2, and f3. While f1 is optional, f2 and f3 are mandatory. Furthermore, f2 consists of either Feature f2.1 or f2.2.
(a) Original Notation [27]
(b) Cardinalities [17]
(c) Constraint
Figure 1: Feature Model Notation
Besides optional and mandatory features, groups of alternative features may have assigned cardinalities [17][18] (or multiplicities [39], respectively). The cardinality of a feature group defines the number of features to be selected from this group. The feature model depicted in Figure 1b represents Software Product 2, which comprises four alternative features (i.e. f4, f5, f6, and f7). The cardinality ⟨2-3⟩ specifies that for Software Product 2 at least two features and at most three features have to be selected.
In [27], typical constraints between features such as requires or mutually exclusive are described in text form. It is now common practice to use a dashed arrow to indicate that a feature requires another one and a dashed double-headed arrow to illustrate that the two marked features must not be co-selected (e.g. [39], [8], and [45]). Finally, Figure 1c shows a feature model that contains a software product with three optional features (i.e. f8, f9, and f10). Due to the constraint requires, when selecting Feature f8, Feature f9 must be selected as well.
2.2 Feature Interactions
In a feature interaction [11][13], in combination multiple features might behave differently than in isolation. Feature interactions have been subject to research in many domains for a long time [11][13][12]. An example of a feature interaction is given by [11] and described in Example 2.2:
{customExample}
[…]
In SPLs, the resolution of feature interaction cannot be hard-coded as during product derivation the features are selected individually obeying the constraints set out by the feature diagram [25]. However, additional code that resolves the unintended behaviour may be conditionally included when two interacting features are present in a software product as proposed for the optional-feature problem. If a feature contains code that depends on another optional feature, the second feature becomes mandatory contrary to its specification, which is known as the optional-feature problem [29]. Different approaches propose extracting the dependent code into a separate module, which is called derivative [35][36] or lifter [38]. Consequently, both features may be included independently and used in software products. However, when using both features in combination, the derivative module is included as well. A derivative module constitutes the resolution of technical dependencies and is not a feature. Consequently, it is not added in the feature model [29]. Furthermore, in order to resolve the interaction of two features they may be marked as mutually exclusive in the feature model, or precedences or priorities may be defined. For a review of resolution techniques for feature interactions we refer interested readers to [48].
2.3 FeatureHouse
FeatureHouse [4] is a framework and tool chain for composing software artefacts to software systems. It uses superimposition to merge the software artefacts, i.e. the substructures of the software artefacts are merged in order to compose the software system. The hierarchical structure of a software artefact is represented by a feature structure tree (FST). For example, a Java artefact may consist of packages, classes, and methods that correspond to FST nodes. Nodes may be non-terminals (i.e. having child nodes) or terminals. Depending on the hierarchy level at which the artefacts shall be merged, the structural elements of the latter need to be chosen as terminal nodes. If, for example, Java artefacts shall be merged at the class level, the classes may be the terminal nodes. Superimposing two FSTs entails merging their nodes beginning with the root node. FeatureHouse supports the composition of artefacts written in various languages, including Java, C#, and XML.
2.4 Feature Refinement
A feature refinement [7] corresponds to a program increment that represents a feature within related software products, e.g. a software product line. Feature refinements comprise software fragments that can be incrementally composed to build a software product whereby the software fragments may be source code and other artefacts. Jak [7] is an extension of the programming language Java that contains key words for feature refinement. A class contained in a Jak file may be refined by another Jak file using the modifier “refines”, e.g. class members and methods may be added.
The AHEAD tools presented in [7], in turn, are able to compose and translate Jak files to Java files. The AHEAD tools also support incrementally composing XML documents written in XAK [3]. XAK is a language to refine XML documents and provides the attribute ”xak:module” to mark a tag as a module (i.e. this tag may be refined) in the base XML file, which is the starting point for composing XML documents. All tags in the XML tree below the module are considered implementation and must not be refined. In separate XAK files, refinements of a module (i.e. XML increment) may be specified. During composition, the refinements are applied to their corresponding module tag in the base XML file, i.e. the tags in the XML-refinement are appended to the corresponding module tag. As XAK refinement files constitute only an increment rather than a complete XML, the files are mostly not schema-compliant.
2.5 Feature Binding
In order to use an optional feature of an SPL, it has to be selected in respective software product. This is called feature binding [16]. Features may be bound at different times. In this context, [27] distinguishes three feature binding times: compile time, start time, and runtime. Features bound at compile time are composed during source code compilation, whereas features bound at start time are selected when a software product is launched. Finally, features bound at runtime may be exchanged while the software product is running. Propositions such as [42] and [41] categorise feature binding times into static binding (i.e. feature binding before software product execution) and dynamic binding (i.e. feature binding during start or while running a software product). Static and dynamic binding each have advantages and disadvantages [10]: On the one hand, dynamic binding entails flexibility with respect to adding and removing features at runtime. On the other hand, this also implies that the software product contains all features. The size of a software product has an impact on the required resources such as memory and CPU. In contrast, static binding requires fewer resources but lacks flexibility at runtime.
Dynamic binding implies adding and removing features during runtime. Hence, it becomes necessary to specify what happens if a running feature shall be removed. [34] summarises the options:
  • The running feature is removed immediately.
  • The current state of the feature is saved before its removal.
  • The running feature is removed after its execution has completed.
  • The running feature is not removed at all.
Furthermore, the removal of a running feature might impact other features with which it shares data or system processes (i.e. background tasks of the operating system) [34]. [34] proposes grouping features into feature binding units at development time that can be bound and unbound at runtime. The feature binding units consider implications of data and system processes the features to be unbound are working on.
A variation point corresponds to the location in an information system at which variability occurs, i.e. features are bound. Methods to implement a variation point (i.e. binding a feature) are called variability mechanisms [26]. In earlier approaches, the binding time of a feature had to be selected at design time by implementing the variability with an appropriate variability mechanism [14][42][10]. For instance, the variability mechanisms discussed in Subsections 2.3 and 2.4 like FeatureHouse [4] and AHEAD [7] allow for static feature binding, solely.
The approach presented in [42] aims to support static and dynamic feature binding without deciding at design time. The code basis of the SPL is developed independently of the feature binding time by using feature refinements. If the features of the SPL shall be bound statically, existing tools are used for composing the feature refinements. In order to bind the features dynamically, the code base is transformed to use the decorator pattern. Features can be selected after compile time by using factory methods of the feature decorators. Thus, all features are included at compile time and available afterwards.
The approach allows selecting either static or dynamic feature binding for all features. In contrast, [41] proposes an approach that builds on the results of [42], but allows selecting the binding time per feature. Changing the order in which features are bound may change their behaviour. Consider Example 2.5.
{customExample}
(Feature Binding Order): Assume Feature A needs to be executed before Feature B. Then, Feature A needs to be bound after Feature B. However, if Feature A is bound statically and Feature B dynamically the order is reversed. The approach presented in [41] ensures the execution order of the features even when features have different binding times. This is achieved by generating hook methods that can be overridden.
In dynamic feature binding, mutually exclusive features are included together in the software product as they are not bound before runtime. Therefore, when binding features at runtime, given constraints (e.g. mutually exclusion and implication) need to be obeyed. Both [42] and [41] use a runtime API that enables checking against the corresponding feature model on whether or not the selected features in an SPL Product are consistent.
Edicts constitute aspects of aspect-oriented programming [14]. For a variation point there may be various edicts. If a feature shall be bound at compile time an edict containing this feature is included. In contrast, if the implementation of the variation point shall be determined at runtime, an edict containing the code of a design pattern (e.g. decorator pattern) that enables choosing a feature at runtime is included.
@@ 2.6 Phases of Software Product Lines Engineering
As opposed to single software products, when developing an SPL the requirements of various similar, yet different, software products need to be collected, managed and implemented. SPL Engineering can be divided into four phases [51][28]:

### v2

@@ Abstract
Different organisations often run similar digitised business processes to achieve their business goals. However, organisations often need to slightly adapt the business processes implemented in an information system in order to adopt them. Various approaches have been proposed to manage variants in process models. While these approaches mainly deal with control flow variability, in previous work we introduced an approach to manage implementation variants of digitised business processes. In this context Software Product Line (SPL) Engineering was applied to manage a set of common core artefacts including a process model from which Process-Aware Information Systems (PAIS) can be derived, which differ in the implementation of their process activities. When deriving a PAIS, implementations are selected for each process activity and then included in the PAIS at compilation time. One challenge that has not yet been solved is giving users of digitised business processes the option of selecting multiple implementations at runtime. This paper extends our previous work by not only allowing for the selection of activity implementations at compile time, but also at start time and runtime. Consequently, it becomes possible to defer the decision as to which implementation should be selected to start time and runtime. Furthermore, multiple implementations of a particular activity may be selected and executed concurrently. The presented approach also allows customising the input and output data of activities. Data from expert interviews with German municipalities suggests digitising business processes with varying implementations is a widespread challenge and our approach is a way to mitigate it.
@@
Business Process Management , Software Product Line Engineering , Process Configuration , Process Variability , Software Reuse , Process Family
[…]
[ innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, ]customExampleExample
[…]
[1]organization=Institute of Databases and Information Systems, Ulm University, addressline=James-Franck-Ring, postcode=89081, city=Ulm, country=Germany
@@ 1 Introduction
The right to self-administration of German municipalities leads to various variants of the same business process. For example, craftspersons may apply for a special parking permit, which allows them to park in areas in which regular citizen have to pay or in which generally parking is not permitted. Special parking permits for craftspersons are available, inter alia, in the German cities of Munich [^1] http://www.muenchen.de/dienstleistungsfinder/muenchen/1072021/, Constance [^2] https://www.konstanz.de/serviceportal/-/leistungen+von+a-z/handwerkerparkausweis-beantragen/vbid6000803, and Stuttgart [^3] https://www.stuttgart.de/vv/leistungen/sonderparkausweise-fuer-gewerbetreibende-und-soziale-dienste.php. The business process for checking an application for the special parking permit is similar between the municipalities. However, there are slight differences among the municipalities. In some municipalities, the special parking permit is issued automatically, whereas in others a municipal employee needs to issue the parking permit manually. Other researchers have observed and investigated business process variability in municipalities as well (La Rosa et al., 2017), focusing on control flow variability.
Approaches have been proposed to deal with business process variability. Reference processes (van der Aalst et al., 2006) and base processes (Hallerbach et al., 2008), respectively, were introduced that may be configured to meet the needs of an individual organisation. However, these approaches focus on variations in the control flow of a business process rather than on the implementation level. Software Product Line (SPL) Engineering (Bass et al., 2013) addresses the challenge of developing and maintaining a set of similar software products. A Software Product Line comprises a common set of core artefacts from which features may be selected in order to build a specific software product. In literature, the process of selecting features from an SPL and building a software product is referred to as product derivation (Deelstra et al., 2005; Kästner and Apel, 2013). A Process-Aware Information System (PAIS) corresponds to a software product that executes a business process involving human actors, applications, and information sources (Dumas et al., 2005). In previous work (Hehnle and Reichert, 2023), we applied and simplified the concepts of SPL Engineering to PAISs (PAIS Product Line). At build time, any implementation can be selected for an activity. However, this static approach lacks flexibility, which is of utmost importance in practice.
@@ 1.1 Problem Statement
Furthermore, the implementations for the activities need to be selected at build time. Selection at start time or runtime is not possible. In order to give the process engineer of a business process the option of selecting features, the decision as to which activity implementations are selected must be deferred to startup time or runtime. In previous work, we used the build tool Apache Maven [^4] https://maven.apache.org/ to conditionally package the selected activity implementations. When activity implementations are to be selected at startup time or runtime a build tool cannot be used anymore. The challenge is to find and adopt a mechanism from SPL Engineering allowing for the selection of activity implementations at startup time and runtime. However, note that when it becomes possible to select and, consequently, deselect activity implementations at runtime, it needs to be specified what shall happen with the data of a running implementation that gets deselected.
@@ 1.2 Contribution
This work addresses the aforementioned limitations and problems of PAIS Product Lines. Building on previous work (Hehnle and Reichert, 2023), the main contribution of this work is threefold:
@@
Moreover, it is assured that the process of designing, developing, selecting activity implementations, and deriving products (known as the phases of SPL Engineering) is tool supported. An evaluation in the form of expert interviews with German municipalities indicates that our contributions tackle a ubiquitous problem and can be applied to a variety of municipal business processes in Germany.
@@ 1.3 Outline
The remainder of this paper is structured as follows: Section 2 presents basic concepts of SPL Engineering, which lay the foundation of the approach, and which are necessary for understanding it. The approach of PAIS Product Lines we introduced in (Hehnle and Reichert, 2023) is described in Section 3. In Section 4, this approach is extended to enable more flexibility in that the data structure may be customised during PAIS Product derivation, multiple implementations for one activity may be selected, and the implementation may be bound at different times. The flexibility of the presented approach is discussed in Section 6. We evaluate our approach by conducting expert interviews with German municipalities in terms of whether the identified problem is widespread and whether our approach mitigates it. Then, we evaluate the validity of our approach. The results are presented in Section 7. Section 8 discusses related work. Section 9 concludes the paper and provides and outlook on open challenges and further research topics.
@@ 2 Fundamentals
In this section, basic concepts of SPL Engineering are presented, which constitute the foundation of our approach. The concepts deal with variability among similar software products that may be derived from a set of common core artefacts. First, the development process (i.e. phases) of an SPL is outlined. Then, according to the phases, it is presented how to collect and sketch the requirements of an SPL as models. Features that represent these requirements can be selected to be included in individual-derived software products, which is called feature binding. Therefore, different feature binding approaches are introduced on an abstract level, before presenting two concrete techniques for feature binding, which are called variability mechanisms. Feature interactions are unintended behaviours which pose a problem when binding different features and are discussed followed by presenting tool support for the entire development process of an SPL (i.e. SPL phases).
@@ 2.1 Phases of Software Product Lines Engineering
As opposed to single software products, when developing an SPL, the requirements of various similar, yet different, software products need to be collected, managed and implemented. SPL Engineering can be divided into four phases (Thüm et al., 2014; Kästner and Apel, 2013):
@@
2.2 Variability Modelling
Kang et al. (Kang et al., 1990) introduced a notation to capture the functionality of a software product from the perspective of a user. By applying this notation, it becomes possible to outline the features of a software product as well as their relationship in a tree-structured feature model. The root node represents the software product. Features of the software product are connected by edges to the root node. Features themselves may consists of other features and are inter-connected with them via edges as well. Little circles at the end of an edge refer to optional features, whereas alternative features are identified by arcs between their corresponding edges. Feature diagrams help developers to identify what has to be configured in a software product.
An example of a feature model is displayed in Figure 1a. Software Product 1 consists of three features, i.e. Features f1, f2, and f3. While f1 is optional, f2 and f3 are mandatory. Furthermore, f2 consists of either Feature f2.1 or f2.2.
(a) Original Notation (Kang et al., 1990)
(b) Cardinalities (Czarnecki et al., 2004)
(c)requires and exclusive Constraint
Figure 1: Feature Model Notation
Besides optional and mandatory features, groups of alternative features may have assigned cardinalities (Czarnecki et al., 2004, 2005) (or multiplicities (Riebisch et al., 2002), respectively). The cardinality of a feature group defines the number of features to be selected from this group. The feature model depicted in Figure 1b represents Software Product 2, which comprises four alternative features (i.e. f4, f5, f6, and f7). The cardinality ⟨2-3⟩ specifies that for Software Product 2 at least two features and at most three features have to be selected.
In (Kang et al., 1990), typical constraints between features such as requires or mutually exclusive are described in text form. It is now common practice to use a dashed arrow to indicate that a feature requires another one and a dashed double-headed arrow to illustrate that the two marked features must not be co-selected (e.g. Riebisch et al. (2002), Benavides et al. (2010), and Schobbens et al. (2007)). Finally, Figure 1c shows a feature model that contains a software product with three optional features (i.e. f8, f9, f10, f11). Due to the constraint requires, when selecting Feature f8, Feature f9 must be selected as well. Furthermore, Features f10 and f11 cannot be co-selected due to the constraint mutually exclusive.
2.3 Feature Binding
In order to use an optional feature of an SPL, it has to be selected in respective software product. This is called feature binding (Czarnecki and Eisenecker, 2000). Features may be bound at different times. In this context, (Kang et al., 1990) distinguishes three feature binding times: compile time, start time, and runtime. Features bound at compile time are composed during source code compilation, whereas features bound at start time are selected when a software product is launched. Finally, features bound at runtime may be exchanged while the software product is running. Propositions such as (Rosenmüller et al., 2008) and (Rosenmüller et al., 2011) categorise feature binding times into static binding (i.e. feature binding before software product execution) and dynamic binding (i.e. feature binding during start or while running a software product). Static and dynamic binding each have advantages and disadvantages (Bosch et al., 2002): On the one hand, dynamic binding entails flexibility with respect to adding and removing features at runtime. On the other hand, this also implies that the software product contains all features. The size of a software product has an impact on the required resources such as memory and CPU. In contrast, static binding requires fewer resources but lacks flexibility at runtime.
Dynamic binding implies adding and removing features during runtime. Hence, it becomes necessary to specify what happens if a running feature shall be removed. (Lee and Kang, 2006) summarises the options:
  • The running feature is removed immediately.
  • The current state of the feature is saved before its removal.
  • The running feature is removed after its execution has completed.
  • The running feature is not removed at all.
Furthermore, the removal of a running feature might impact other features with which it shares data or system processes (i.e. background tasks of the operating system) (Lee and Kang, 2006). Lee and Kang (2006) propose grouping features into feature binding units at development time that can be bound and unbound at runtime. The feature binding units consider implications of data and system processes the features to be unbound are working on.
A variation point corresponds to the location in an information system at which variability occurs, i.e. features are bound. Methods to implement a variation point (i.e. binding a feature) are called variability mechanisms (Jacobson et al., 1997). In earlier approaches, the binding time of a feature had to be selected at design time by implementing the variability with an appropriate variability mechanism (Chakravarthy et al., 2008; Rosenmüller et al., 2008; Bosch et al., 2002).
The approach presented in (Rosenmüller et al., 2008) aims to support static and dynamic feature binding without deciding at design time. The code basis of the SPL is developed independently of the feature binding time by using feature refinements. If the features of the SPL shall be bound statically, existing tools are used for composing the feature refinements. In order to bind the features dynamically, the code base is transformed to use the decorator pattern (also known as wrapper (Gamma, 1995)). Features can be selected after compile time by using factory methods of the feature decorators. Thus, all features are included at compile time and available afterwards.
The approach allows selecting either static or dynamic feature binding for all features. In contrast, Rosenmüller et al. (2011) proposes an approach that builds on the results of Rosenmüller et al. (2008) but allows selecting the binding time per feature. Changing the order in which features are bound may change their behaviour. Consider Example 2.3.
{customExample}
(Feature Binding Order): Assume Feature A needs to be executed before Feature B. Then, Feature A needs to be bound after Feature B. However, if Feature A is bound statically and Feature B dynamically the order is reversed. The approach presented in (Rosenmüller et al., 2011) ensures the execution order of the features even when features have different binding times. This is achieved by generating hook methods that can be overridden.
In dynamic feature binding, mutually exclusive features are included together in the software product as they are not bound before runtime. Therefore, when binding features at runtime, given constraints (e.g. mutually exclusion and implication) need to be obeyed. Both Rosenmüller et al. (2008) and Rosenmüller et al. (2011) use a runtime API that enables checking against the corresponding feature model on whether or not the selected features in an SPL Product are consistent.
Edicts constitute aspects of aspect-oriented programming (Chakravarthy et al., 2008). For a variation point there may be various edicts. If a feature shall be bound at compile time an edict containing this feature is included. In contrast, if the implementation of the variation point shall be determined at runtime, an edict containing the code of a design pattern (e.g. decorator pattern) that enables choosing a feature at runtime is included.
2.4 FeatureHouse
FeatureHouse (Apel et al., 2009) is variability mechanism for static feature binding and a framework and tool chain for composing software artefacts to software systems. It uses superimposition to merge the software artefacts, i.e. the substructures of the software artefacts are merged in order to compose the software system. The hierarchical structure of a software artefact is represented by a feature structure tree (FST). For example, a Java artefact may consist of packages, classes, and methods that correspond to FST nodes. Nodes may be non-terminals (i.e. having child nodes) or terminals. Depending on the hierarchy level at which the artefacts shall be merged, the structural elements of the latter need to be chosen as terminal nodes. If, for example, Java artefacts shall be merged at the class level, the classes may be the terminal nodes. Superimposing two FSTs entails merging their nodes beginning with the root node. FeatureHouse supports the composition of artefacts written in various languages, including Java, C#, and XML.
2.5 Feature Refinement
A feature refinement (Batory et al., 2004) corresponds to a program increment that represents a feature within related software products, e.g. a software product line. Feature refinements comprise software fragments that can be incrementally composed to build a software product whereby the software fragments may be source code and other artefacts. Jak (Batory et al., 2004) is an extension of the programming language Java that contains key words for feature refinement. A class contained in a Jak file may be refined by another Jak file using the modifier “refines”, e.g. class members and methods may be added.
The AHEAD tools presented in (Batory et al., 2004), in turn, are able to compose and translate Jak files to Java files, which constitutes another variability mechanism for static feature binding. The AHEAD tools also support incrementally composing XML documents written in XAK (Anfurrutia et al., 2007). XAK is a language to refine XML documents and provides the attribute ”xak:module” to mark a tag as a module (i.e. this tag may be refined) in the base XML file, which is the starting point for composing XML documents. All tags in the XML tree below the module are considered implementation and must not be refined. In separate XAK files, refinements of a module (i.e. XML increment) may be specified. During composition, the refinements are applied to their corresponding module tag in the base XML file, i.e. the tags in the XML-refinement are appended to the corresponding module tag. As XAK refinement files constitute only an increment rather than a complete XML, the files are mostly not schema-compliant.
2.6 Feature Interactions
In a feature interaction (Bowen T.F. et al., 1989; Cameron and Velthuijsen, 1993), in combination multiple features might behave differently than in isolation. Feature interactions have been subject to research in many domains for a long time (Bowen T.F. et al., 1989; Cameron and Velthuijsen, 1993; Calder et al., 2003). An example of a feature interaction is given by (Bowen T.F. et al., 1989) and described in Example 2.6:
{customExample}
[…]
In SPLs, the resolution of feature interaction cannot be hard-coded as during product derivation the features are selected individually obeying the constraints set out by the feature diagram (Hunt and McGregor, 2007). However, additional code that resolves the unintended behaviour may be conditionally included when two interacting features are present in a software product as proposed for the optional-feature problem. If a feature contains code that depends on another optional feature, the second feature becomes mandatory contrary to its specification, which is known as the optional-feature problem (Kästner et al., 2009). Different approaches propose extracting the dependent code into a separate module, which is called derivative (Liu, 2005; Liu et al., 2006) or lifter (Prehofer, 1997). Consequently, both features may be included independently and used in software products. However, when using both features in combination, the derivative module is included as well. A derivative module constitutes the resolution of technical dependencies and is not a feature. Consequently, it is not added in the feature model (Kästner et al., 2009). Furthermore, in order to resolve the interaction of two features they may be marked as mutually exclusive in the feature model, or precedences or priorities may be defined. For a review of resolution techniques for feature interactions we refer interested readers to (Soares et al., 2018).